OpenStudy (anonymous):

What is the solution to the equation? 5(x + 4) - 4x = 18 - 23 x=?

OpenStudy (adajiamcneal):

if you go on cym@th.com they will give you all the work and the answer :) you just have to type in the equation

OpenStudy (anonymous):

5x + 20 - 4x = -5 x + 20 =-5 - 20 -20 x = -25

OpenStudy (anonymous):

thanks @adajiamcneal this helps a lot :)))

OpenStudy (adajiamcneal):

your welcome :)